The horse on a carousel is 4.5 m from the central axis.

a. If the carousel rotates at 0.15 rev/s , how long does it take the horse to go around twice?

b. How fast is a child on the horse going (in m/s)?
Answer #1

(a) Time taken by carousel to go around twice = 2/0.15 = 13.3 s

(b) Speed os child on horse = 2×π× 4.5× 0.15 m/s = 4.24 m/s